OpenStudy (anonymous):

what is 0.69 (repeating line over the 69) expressed as the quotient of two integers in simplest form??

OpenStudy (misty1212):

HI!!

DivineSolar (divinesolar):

Im here!

OpenStudy (misty1212):

the real quick answer is \[\frac{69}{99}\] which reduces

DivineSolar (divinesolar):

Misty i have this you can handle more important things :). Besides you need an explanation for that.

DivineSolar (divinesolar):

Or else that is breaking the code of conduct

OpenStudy (misty1212):

the method is \[x=.696969...\\ 100x=69.696969...\] subtract the first from the second and get \[99x=69\] making \[x=\frac{69}{99}\]

DivineSolar (divinesolar):

Hehe my explanation would have been simpler but okay.

OpenStudy (misty1212):

method two is learning how to sum a geometric series

OpenStudy (misty1212):

love to see it

DivineSolar (divinesolar):

And by the way it would be over 100 :)

DivineSolar (divinesolar):

not 99

OpenStudy (misty1212):

ok now you are wrong

OpenStudy (misty1212):

it is not \(.69\) it is \(\overline{.69}\)

DivineSolar (divinesolar):

Oh yeah sorry 99 :)

DivineSolar (divinesolar):

Forgot about the repeating line >.<

OpenStudy (misty1212):

sure i got time for one more then i have to run

OpenStudy (anonymous):

Which number is irrational? A. \[\frac{ 1}{ 6 }\] B. \[\sqrt{17}\] C. \[\sqrt{49}\] D. 0.06 (repeated)